Danielle's test scores are 86,80,84, and 90. The final exam will count as two-thirds of the final grade. What score does Danielle need on the final in order to earn a B, which requires an average score of 80? What score does she need to earn an A, which requires an everage of 90?
Answer by stanbon(75887)

Danielle's test scores are 86,80,84, and 90. The final exam will count as two-thirds of the final grade. What score does Danielle need on the final in order to earn a B, which requires an average score of 80? What score does she need to earn an A, which requires an everage of 90?
-----
Of the total 100 points offered,
33 1/3 points are available based on the 1st 4 tests::
She earned [(86+80+84+90)/4] = 85% of those points = 0.85*33 1/3 = 28 1/3
------
66 2/3 points are available based on the final test::
For a B she needs 80-28 1/3 = 51 2/3 points
---
She needs 51 2/3 of the 66 2/3 = 51.667/66.667 = 77.5% on the final to get a B
-----------------------
She needs 61 2/3 of the 66 2/3 points or 61.667/66.667 = 92.5% on
the final to get an A
----------
Cheers,
Stan H.